You dilute a natural sample 10-4, plate 100 microliters, and observe 200 colonies on the plate. what was the initial concentration, in units of cells/ml?
Answer #1

100 microliters = 0.1 ml

Number of cells observed = 200

Dilution = 10-4

Initial concentration of cells = 200 x 1/10

                                          = 200 / 10 = 20 cells/ml